# Evaluating Memory in LLM Agents via Incremental Multi-Turn Interactions

### v1

@@ Abstract
Recent benchmarks for Large Language Model (LLM) agents primarily focus on evaluating reasoning, planning, and execution capabilities, while another critical component—memory, encompassing how agents memorize, update, and retrieve long-term information—is under-evaluated due to the lack of benchmarks. We term agents with memory mechanisms as memory agents. In this paper, we identify four core competencies essential for memory agents: accurate retrieval, test-time learning, long-range understanding, and conflict resolution. Existing datasets either rely on limited context lengths or are tailored for static, long-context settings like book-based QA, which do not reflect the interactive, multi-turn nature of memory agents that incrementally accumulate information. Furthermore, no existing benchmarks cover all four competencies. Therefore, we introduce MemoryAgentBench, a new benchmark specifically designed for memory agents. Our benchmark combines reformulated existing datasets with newly constructed ones, covering the above four memory competencies, providing a systematic and challenging testbed for assessing memory quality. We evaluate a diverse set of memory agents, ranging from simple context-based and retrieval-augmented generation (RAG) systems to advanced agents with external memory modules and tool integration. Empirical results reveal that current methods fall short of mastering all four competencies, underscoring the need for further research into comprehensive memory mechanisms for LLM agents.
@@ 1 Introduction
Large Language Model (LLM) agents have rapidly transitioned from proof-of-concept chatbots to end-to-end systems that can write software [42], control browsers [34], and reason over multi-modal inputs. Frameworks such as Manus, OWL [16], OpenHands [42], and Codex routinely solve complex, tool-rich tasks and achieve state-of-the-art results on agentic benchmarks like GAIA [31] and SWE-Bench [18]. Yet these evaluations focus almost exclusively on reasoning (planning, tool using, code synthesis) and leave the equally important question of memorization (abstraction, storing, updating, retrieving) largely under-explored.
@@
Recent memory-centric architectures—ranging from parametric memory systems like MemoryLLM [44], SELF-PARAM [43], and M+[46] to commercial token-level memory solutions such as MemGPT[36, 27], Mem0[7], Cognee, and Zep[37]—employ diverse strategies for storing and retrieving past information. Despite growing interest, their real-world effectiveness remains largely anecdotal, and there is currently no unified benchmark for systematically evaluating the quality of memory in agents. In this paper, we refer to agents equipped with memory mechanisms as Memory Agents, where memory can take various forms, including parameters, vectors, textual histories, or external databases. In this paper, we primarily focus on memory agents that utilize textual histories and external databases, as these approaches are most commonly deployed in real-world applications. In contrast, memory encoded in model parameters [44, 46, 51] remains largely within academic research and is typically less capable than proprietary memory systems equipped on closed-sourced API models.
@@
To evaluate memory agents, we identify four complementary competencies (Examples shown in Figure 1): (1) Accurate Retrieval: The ability to extract the correct snippet in response to a query. This can involve one-hop or multi-hop retrieval, as long as the relevant information can be accessed with a single query. (2) Test-Time Learning: The capacity to incorporate new behaviors or acquire new skills during deployment, without additional training. (3) Long-Range Understanding: The ability to integrate information distributed across extended contexts ( 100k tokens) and answer questions requiring a global understanding of the entire sequence. (4) Conflict Resolution: The skill to revise, overwrite, or remove previously stored information when faced with contradictory evidence, aligning with goals in model editing and knowledge unlearning tasks [30, 45].
@@
Previous datasets developed to evaluate memory in language models have notable limitations. Early benchmarks such as LOCOMO [29] ( 9k tokens), LooGLE[23] ( 24k tokens), and LongBench[3] ( 20k tokens) feature relatively short contexts that no longer challenge current models. More recent datasets like NovelQA[40] (200k tokens), NOCHA[19] (127k tokens), Loong[41] (100k tokens), and -Bench[53] (150k tokens) extend the context length to evaluate global reasoning and retrieval capabilities. However, these datasets were primarily designed for evaluating long-context language models rather than memory agents. The reason that long-context benchmarks cannot be directly used to evaluate memory agents is as follows. There is a fundamental distinction between memory and long context: memory serves as a compressed and distilled representation of past information. Rather than storing all historical content verbatim, memory selectively extracts salient details, removes irrelevant information, and often incorporates new inferences derived from prior experiences. Consequently, memory agents are designed to process context incrementally—absorbing input piece by piece, abstracting and consolidating information over time, generating new inferences, and learning novel rules from accumulated history. For this reason, datasets that provide the entire context in a single block are not directly applicable to evaluating memory agents. A more recent effort, LongMemEval [48], seeks to address this limitation by using synthetic long-form conversations, which can be injected into memory gradually, session by session. Nonetheless, its evaluation framework remains constrained by limited topical diversity and less realistic interaction patterns, reducing its applicability to real-world memory agent scenarios.
@@
To address these limitations, we introduce a unified benchmark framework, MemoryAgentBench, specifically designed to evaluate a broad spectrum of memory mechanisms in agent systems. We also provide a framework for memory agent evaluation. In this framework, agents are presented with sequences of textual inputs that simulate multi-turn interactions with users. We repurpose existing datasets originally developed for long-context LLM evaluation by segmenting their inputs into multiple chunks and feeding them incrementally to the agent. However, since these datasets do not fully capture all four targeted memory competencies, we also introduce two new datasets: EventQA and FactConsolidation, designed to evaluate accurate retrieval and conflict resolution, respectively. Our benchmark includes evaluations of state-of-the-art commercial memory agents (such as Mem0 and MemGPT), long-context agents that treat the full input as memory, and RAG agents that extend their memory through retrieval methods. We examine how techniques developed for long-context models and RAG transfer to the memory agent setting, and how commercial memory agents perform under more challenging, competency-specific tests. By providing a consistent evaluation protocol across diverse agent architectures and datasets, MemoryAgentBench delivers comprehensive insights into agent performance across the four core memory competencies.
@@
Datasets: We re-structure existing datasets and create two new datasets to construct a comprehensive benchmark, covering four distinct memory competencies.
@@ 2.1 Benchmarks with Long Input
In this section, we review prior work on long-context benchmarks. Early benchmarks designed for long-context evaluation include LongBench[3] and LooGLE[23], with average input lengths of approximately 20k and 24k tokens, respectively. More recent benchmarks—such as -Bench [53], HELMET[50], RULER[15], NOCHA[19], NoLiMa [33] and LongBench V2[4]—extend context lengths to over 100k tokens and are primarily intended to evaluate the capabilities of long-context models. However, despite their scale, these benchmarks are not designed to assess memory agents, and no prior work has repurposed them for that goal. More recently, LOCOMO [29] and LongMemEval [48] have been proposed specifically for evaluating memory agents. While promising, LOCOMO still features relatively short conversations (9k), and LongMemEval uses synthetic conversations with limited topical diversity, making the dialogues less realistic and potentially less representative of real-world memory use cases.
@@ 2.2 Agents with Memory Mechanisms
Memory mechanisms are attracting more and more attention lately [47]. Recent advancements in LLMs have demonstrated the capability to process extended context lengths, ranging from 100K to over 1 million tokens. For instance, models such as GPT-4o [35] and Claude 3.7 [1] can handle inputs of approximately 100K to 200K tokens, while models like Gemini 2.0 Pro [8] and the GPT-4.1 series extend this capacity beyond 1 million tokens. These strong long-context capabilities enable a simple yet effective form of memory: storing information directly within the context window. However, this approach is inherently constrained by a hard limit—once the context window is exceeded, earlier information must be discarded.
@@
In parallel, RAG continues to serve as a dominant paradigm for managing excessive context. By retrieving relevant information from earlier context and feeding it to the LLM, RAG allows systems to overcome context length limitations. For example, OpenAI’s recent memory functionality111https://openai.com/index/memory-and-new-controls-for-chatgpt/ combines explicit user preference tracking with retrieval-based methods that reference prior interactions. RAG methods can be broadly classified into three categories: 1. Simple RAG: These methods rely on string-matching techniques such as TF-IDF, BM25 [38], and BMX [25], which are entirely non-neural and operate on string-level similarity. 2. Embedding-based RAG: This class leverages neural encoders, primarily transformers, to map text into dense vector representations [49]. Early methods like DPR [20] and Contriever [17] are based on BERT [9], while more recent models such as NV-Embed-v2 [22] utilize decoder-only backbones and achieve significantly improved retrieval performance. 3. Structure-Augmented RAG: These approaches enhance retrieval with structural representations such as graphs or trees. Representative systems include GraphRAG [10], RAPTOR [39], HippoRAG-V2 [12], Cognee, Zep [37], and Mem0 [7], where Mem0 also offers a graph-augmented variant, Mem0g, built on structured factual knowledge. Despite their effectiveness, RAG-based methods face challenges with ambiguous queries, multi-hop reasoning, and long-range comprehension. When questions require integrating knowledge across an entire session or learning from long, skill-encoding inputs, the retrieval mechanism—limited to the top-k most relevant passages—may fail to surface the necessary information. To address these limitations, Agentic Memory Agents introduce an iterative, decision-driven framework. Rather than relying on a single-pass retrieval, these agents dynamically process the query, retrieve evidence, reflect, and iterate through multiple retrieval and reasoning cycles. Examples include MemGPT [36], Self-RAG [2], and Auto-RAG [52]. This agentic design is particularly effective for resolving ambiguous or multi-step queries. Nonetheless, these methods remain fundamentally constrained by the limitations of RAG—namely, the inability to fully understand or learn from long-range context that is inaccessible via retrieval alone.
@@ 3 MemoryAgentBench
3.1 Aspects of the Evaluation
The evaluation of memory agents encompasses the following key dimensions:
Accurate Retrieval (AR)
The task of accurately retrieving information has been extensively explored in prior work. In the domain of long-context modeling, the Needle-in-a-Haystack (NIAH) task is widely used to evaluate a model’s ability to locate a specific value based on a given key within a lengthy input. Extensions such as multi-value NIAH further test the model’s capacity to retrieve multiple values scattered across the input context. In the RAG setting, this corresponds to document-based QA, where the model must identify and extract relevant snippets from one or more documents to answer a query. These snippets might reside in a single location or be distributed across multiple documents. In this paper, we focus on agentic settings, where the “long context” or “multiple documents” become long-form conversations. We define Accurate Retrieval (AR) as the ability of an agent to identify and retrieve important information that may be dispersed throughout a long dialogue history.
Test-Time Learning (TTL)
Long-Range Understanding (LRU)
Conflict Resolution (CR)
In long-term interactions, agents often face evolving or conflicting information—whether about the external world (e.g., changes in political leadership) or user-specific facts (e.g., a new occupation). This challenge is closely related to model editing [30, 11] and knowledge unlearning [45], which focus on modifying or removing factual knowledge from language models. We define Conflict Resolution (CR) as the agent’s ability to detect and resolve contradictions between existing knowledge and newly acquired information, ensuring the agent remains aligned with current realities and user states. CR is distinct from Abstractive Retrieval (AR) in two key ways. (1) Certain questions requiring CR cannot be answered solely through AR. As illustrated in Figure 1, an agent that retrieves all facts related to pears may fail to identify the updated information in the second message. (2) In AR, earlier messages remain relevant and should be retained, even when multiple pieces of evidence are required. In contrast, CR involves identifying outdated or incorrect information and discarding it. That is, AR requires preservation of all related content, whereas CR requires overwriting prior facts to reflect the most up-to-date truth.
| Capability | Benchmarks / Tasks | # of Sequences | # of QAs | Avg Len |
| Accurate Retrieval | RULER-QA [15] | 2 | 200 | 309K |
| RULER-NIAH-MQ [15] | 1 | 100 | 448K | |
| Bench-QA [53] | 100 | 100 | 183K | |
| LongMemEval (S*) [48] | 5 | 300 | 355K | |
| EventQA (ours) | 5 | 500 | 534K | |
| Test-Time Learning | BANKING-77 | 1 | 100 | |
| CLINC-150 | 1 | 100 | ||
| NLU | 1 | 100 | 103K | |
| TREC (Coarse) | 1 | 100 | ||
| TREC (Fine) | 1 | 100 | ||
| Movie-Rec Redial [14] | 1 | 200 | 1.44M | |
| Long-Range Understanding | Bench-Sum [53] | 100 | 100 | 172K |
| Conflict Resolution | FactConsolidation-SH (ours) | 1 | 100 | 262K |
| FactConsolidation-MH (ours) | 1 | 100 |
@@ 3.2 Dataset Preperation
In this section, we describe how we adopt existing datasets and construct new ones for evaluating each aspect introduced in Section 3.1. All datasets with their categories are shown in Table 1. We introduce the details in datasets curation in Appendix A.
@@
We adopt five datasets to evaluate the accurate retrieval capability of memory agents. Four are adapted from existing benchmarks, and one is newly constructed: (1) RULER-QA: This is a NIAH-style QA task where a long passage contains single (QA-1) or multiple (QA-2) snippets answering the input question. The agent must identify and extract relevant snippets from the extended context. (2) NIAH-MQ: We use the multiple-query (MQ) version of the NIAH dataset from RULER [15], where each query seeks a different numeric value embedded in a long passage. The agent must accurately retrieve multiple distinct answers. (3) Bench-En.QA: This task from Bench presents free-form QA questions based on entire books, with all entities replaced by fictitious names to avoid contamination from model pretraining. Compared to synthetic datasets like RULER-QA, this benchmark is more realistic and challenging due to the natural narrative structure of books. (4) LongMemEval: This benchmark evaluates memory agents on long dialogue histories. Although task types like information extraction (IE) or multi-session reasoning are included, most tasks can be reformulated as single-retrieval problems requiring agents to retrieve the correct segments spanning a long multi-turn conversation. Among these, LongMemEval is already formatted for agent-based evaluation with session separation. We use the original LongMemEval(S) dataset (110K tokens) and reformulated chat history into five long dialogues (355K tokens) with 300 questions (LongMemEval (S*) in Table 1). We create LongMemEval (S*) specifically for increasing the number of questions per context, mitigating the exhaustive needs of reconstructing the memory for each quesiton. (5) EventQA (ours): We introduce EventQA this reasoning style NIAH task to evaluate agents’ ability to recall and reason about temporal sequences in long-form narratives. In this dataset, the agent is required to read a novel and select the correct event from a series of candidates after receiving up-to five previous events. For these datasets, which are originally designed for long-context modeling, we split documents into chunks and sequentially inject them into the agent.
@@
We evaluate TTL via two task categories: (1) Multi-Class Classification (MCC): We adopt five classification datasets used in prior TTL work [5, 50]: BANKING77 [6], CLINC150 [21], TREC-Coarse, TREC-Fine [26], and NLU [28]. Each task requires the agent to map sentences to class labels, leveraging previously seen labeled examples in context. (2) Recommendation (Recom): We use the Redial [24] dataset to evaluate movie recommendation via dialogue. Following the setup from He et al. [13], the agent is exposed to thousands of movie-related dialogue turns and is asked to recommend twenty relevant movies based on the long interaction history.
@@
We adopt the Summarization task En.Sum from -Bench [53]. The agent is required to analyze and organize the plot and characters of the novel, and then compose a summary of 1000 to 1200 words.
Datasets for Conflict Resolution (CR)
To assess whether an agent can consolidate conflicting factual updates and reason over them, we construct a new dataset called FactConsolidation. Specifically, We build this benchmark using counterfactual edit pairs from MQUAKE [54]. Each pair contains a true fact and a rewritten, contradictory version. These are ordered such that the rewritten (new) fact appears after the original, simulating a realistic update scenario. We concatenate multiple such edit pairs to create long contexts of length 32K, 64K, 262K. We then adpot MQUAKE’s original questions and categorize them into: (1) FactConsolidation-SH (Ours) (SH means Single-Hop), requiring direct factual recall (e.g., “Which country was tool created in?”), and (2) FactConsolidation-MH (Ours) (MH refers to Multi-Hop), requiring inference over multiple facts (e.g., “What is the location of death of the spouse of person ?”). Agents are prompted to prioritize later information in case of conflict and reason based on the final memory state. This setup directly evaluates the strength and consistency of conflict resolution over long sequences.

### v2

@@ Abstract
Recent benchmarks for Large Language Model (LLM) agents primarily focus on evaluating reasoning, planning, and execution capabilities, while another critical component—memory, encompassing how agents memorize, update, and retrieve long-term information—is under-evaluated due to the lack of benchmarks. We term agents with memory mechanisms as memory agents. In this paper, based on classic theories from memory science and cognitive science, we identify four core competencies essential for memory agents: accurate retrieval, test-time learning, long-range understanding, and selective forgetting. Existing benchmarks either rely on limited context lengths or are tailored for static, long-context settings like book-based QA, which do not reflect the interactive, multi-turn nature of memory agents that incrementally accumulate information. Moreover, no existing benchmarks cover all four competencies. We introduce MemoryAgentBench, a new benchmark specifically designed for memory agents. Our benchmark transforms existing long-context datasets and incorporates newly constructed datasets into a multi-turn format, effectively simulating the incremental information processing characteristic of memory agents. By carefully selecting and curating datasets, our benchmark provides comprehensive coverage of the four core memory competencies outlined above, thereby offering a systematic and challenging testbed for assessing memory quality. We evaluate a diverse set of memory agents, ranging from simple context-based and retrieval-augmented generation (RAG) systems to advanced agents with external memory modules and tool integration. Empirical results reveal that current methods fall short of mastering all four competencies, underscoring the need for further research into comprehensive memory mechanisms for LLM agents.
@@ 1 Introduction
Large Language Model (LLM) agents have rapidly transitioned from proof-of-concept chatbots to end-to-end systems that can write software (Wang et al., 2024c), control browsers (Müller & Žunič, 2024), and reason over multi-modal inputs. Frameworks such as Manus, OWL (Hu et al., 2025), OpenHands (Wang et al., 2024c), and Codex routinely solve complex, tool-rich tasks and achieve state-of-the-art results on agentic benchmarks like GAIA (Mialon et al., 2023) and SWE-Bench (Jimenez et al., 2023). Yet these evaluations focus almost exclusively on reasoning (planning, tool using, code synthesis) and leave the equally important question of memorization (abstraction, storing, updating, retrieving) largely under-explored. Recent memory-centric architectures—ranging from parametric memory systems like MemoryLLM (Wang et al., 2024d), SELF-PARAM (Wang et al., ), and M+(Wang et al., 2025) to commercial token-level memory solutions such as MemGPT(Packer et al., 2023; Lin et al., 2025), Mem0(Chhikara et al., 2025), Cognee(Markovic et al., 2025), and Zep(Rasmussen et al., 2025)—employ diverse strategies for storing and retrieving past information. Despite growing interest, their real-world effectiveness remains largely anecdotal, and there is currently no unified benchmark for systematically evaluating the quality of memory in agents. In this paper, we refer to agents equipped with memory mechanisms as Memory Agents, where memory can take various forms, including parameters, vectors, textual histories, or external databases. In this paper, we primarily focus on memory agents that utilize textual histories and external databases, as these approaches are most commonly deployed in real-world applications. In contrast, memory encoded in model parameters (Wang et al., 2024d; 2025; Yin et al., 2024) remains largely within academic research and is typically less capable than proprietary memory systems equipped on closed-sourced API models.
@@
Based on some classic theories in memory and cognitive science (James, 1890; McClelland et al., 1995; Anderson & Neely, 1996; Wimber et al., 2015), we identify four complementary competencies (Examples shown in Figure 1) to evaluate memory agents: (1) Accurate Retrieval (AR): The ability to extract the correct snippet in response to a query. This can involve one-hop or multi-hop retrieval, as long as the relevant information can be accessed with a single query. (2) Test-Time Learning (TTL): The capacity to incorporate new behaviors or acquire new skills during deployment, without additional training. (3) Long-Range Understanding (LRU): The ability to integrate information distributed across extended contexts ( 100k tokens) and answer questions requiring a global understanding of the entire sequence. (4) Selective Forgetting (SF): The skill to revise, overwrite, or remove previously stored information when faced with contradictory evidence, aligning with goals in model editing and knowledge unlearning tasks (Meng et al., 2023; Wang et al., 2024e). For these fuor competencies, we provide more detailed definitions in Appendix A.
@@
Previous datasets developed to evaluate memory in language models have notable limitations. Early benchmarks such as LOCOMO (Maharana et al., 2024) ( 9k tokens), LooGLE(Li et al., 2023) ( 24k tokens), and LongBench(Bai et al., 2023) ( 20k tokens) feature relatively short contexts that no longer challenge current models. More recent datasets like NovelQA(Wang et al., 2024a) (200k tokens), NOCHA(Karpinska et al., 2024) (127k tokens), Loong(Wang et al., 2024b) (100k tokens), and -Bench(Zhang et al., 2024) (150k tokens) extend the context length to evaluate global reasoning and retrieval capabilities. However, these datasets were primarily designed for evaluating long-context language models rather than memory agents. The reason that long-context benchmarks cannot be directly used to evaluate memory agents is as follows. There is a fundamental distinction between memory and long context: memory serves as a compressed and distilled representation of past information. Rather than storing all historical content verbatim, memory selectively extracts salient details, removes irrelevant information, and often incorporates new inferences derived from prior experiences. Consequently, memory agents are designed to process context incrementally—absorbing input piece by piece, abstracting and consolidating information over time, generating new inferences, and learning novel rules from accumulated history. For this reason, datasets that provide the entire context in a single block are not directly applicable to evaluating memory agents. A more recent effort, LongMemEval (Wu et al., 2025), seeks to address this limitation by using synthetic long-form conversations, which can be injected into memory gradually, session by session. Nonetheless, its evaluation framework remains constrained by limited topical diversity and less realistic interaction patterns, reducing its applicability to real-world memory agent scenarios.
@@
To address these limitations, we introduce a unified benchmark framework, MemoryAgentBench, specifically designed to evaluate a broad spectrum of memory mechanisms in agent systems. We also provide a framework for memory agent evaluation. In this framework, agents are presented with sequences of textual inputs that simulate multi-turn interactions with users. We reconstructed existing datasets originally developed for long-context LLM evaluation by segmenting and reconstructing inputs into multiple dialogue chunks and feeding them incrementally to the agent in a time order. However, since these datasets do not fully capture all four targeted memory competencies, we also introduce two new datasets: EventQA and FactConsolidation, designed to evaluate accurate retrieval and selective forgetting, respectively. Our benchmark includes evaluations of state-of-the-art commercial memory agents (such as MIRIX and MemGPT), long-context agents that treat the full input as memory, and RAG agents that extend their memory through retrieval methods. We examine how techniques developed for long-context models and RAG transfer to the memory agent setting. By providing a consistent evaluation protocol across diverse agent architectures and datasets, MemoryAgentBench delivers comprehensive insights into agent performance across the four core memory competencies.
@@
Datasets: We reconstruct existing datasets and create two new datasets to construct a comprehensive benchmark, covering four distinct memory competencies.
@@ 2.1 Benchmarks with Long Input
In this section, we review prior work on long-context benchmarks. Early benchmarks designed for long-context evaluation include LongBench(Bai et al., 2023) and LooGLE(Li et al., 2023), with average input lengths of approximately 20k and 24k tokens, respectively. More recent benchmarks—such as -Bench (Zhang et al., 2024), HELMET(Yen et al., 2024), RULER(Hsieh et al., 2024), NOCHA(Karpinska et al., 2024), NoLiMa (Modarressi et al., 2025) and LongBench V2(Bai et al., 2024)—extend context lengths to over 100k tokens and are primarily intended to evaluate the capabilities of long-context models. However, despite their scale, these benchmarks are not designed to assess memory agents, and no prior work has repurposed them for that goal. More recently, LOCOMO (Maharana et al., 2024), LongMemEval (Wu et al., 2025), RealTalk (Lee et al., 2025) and StoryBench (Wan & Ma, 2025) have been proposed specifically for evaluating memory agents. While promising, LOCOMO still features relatively short conversations (9k), and LongMemEval uses synthetic conversations with limited topical diversity, making the dialogues less realistic and potentially less representative of real-world memory use cases. Meanwhile, the evaluation scope of the above benchmarks is not sufficient to comprehensively assess long-term memory from multiple dimensions.
@@ 2.2 Agents with Memory Mechanisms
Memory mechanisms are attracting more and more attention lately (Wang et al., 2025/02). Recent advancements in LLMs have demonstrated the capability to process extended context lengths, ranging from 100K to over 1 million tokens. For instance, models such as GPT-4o (OpenAI, 2025b) and Claude 3.7 (Anthropic, 2025) can handle inputs of approximately 100K to 200K tokens, while models like Gemini 2.0 Pro (DeepMind, 2025) and the GPT-4.1 series extend this capacity beyond 1 million tokens. These strong long-context capabilities enable a simple yet effective form of memory: storing information directly within the context window. However, this approach is inherently constrained by a hard limit—once the context window is exceeded, earlier information must be discarded.
@@
In parallel, RAG continues to serve as a dominant paradigm for managing excessive context. By retrieving relevant information from earlier context and feeding it to the LLM, RAG allows systems to overcome context length limitations. For example, OpenAI’s recent memory functionality111https://openai.com/index/memory-and-new-controls-for-chatgpt/ combines explicit user preference tracking with retrieval-based methods that reference prior interactions. RAG methods can be broadly classified into three categories: Simple RAG: These methods rely on string-matching techniques such as TF-IDF, BM25 (Robertson & Walker, 1994), and BMX (Li et al., 2024), which are entirely non-neural and operate on string-level similarity. Embedding-based RAG: This class leverages neural encoders, primarily transformers, to map text into dense vector representations (Wu et al., 2022). Early methods like DPR (Karpukhin et al., 2020) and Contriever (Izacard et al., 2021) are based on BERT (Devlin et al., 2019), while more recent models such as Qwen3-Embedding (Zhang et al., 2025) achieve significantly improved retrieval performance. Structure-Augmented RAG: These approaches enhance retrieval with structural representations such as graphs or trees. Representative systems include GraphRAG (Edge et al., 2024), RAPTOR (Sarthi et al., 2024), HippoRAG-V2 (Gutiérrez et al., 2025), Cognee, Zep (Rasmussen et al., 2025), MemoRAG (Qian et al., 2025), Mem0 (Chhikara et al., 2025), MemoryOS (Kang et al., 2025), Memary (kingjulio8238 & Memary contributors, 2024) and Memobase (memodb-io & Memobase contributors, 2025). Despite their effectiveness, RAG-based methods face challenges with ambiguous queries, multi-hop reasoning, and long-range comprehension. When questions require integrating knowledge across an entire session or learning from long, skill-encoding inputs, the retrieval mechanism—limited to the top-k most relevant passages—may fail to surface the necessary information. To address these limitations, Agentic Memory Agents introduce an iterative, decision-driven framework. Rather than relying on a single-pass retrieval, these agents dynamically process the query, retrieve evidence, reflect, and iterate through multiple retrieval and reasoning cycles. Examples include MemGPT (Packer et al., 2023), Self-RAG (Asai et al., 2023), Auto-RAG (Yu et al., 2024), A-MEM (Xu et al., 2025), Mem1 (Zhou et al., 2025), MemAgent (Yu et al., 2025), and MIRIX (Wang & Chen, 2025). This agentic design is particularly effective for resolving ambiguous or multi-step queries. Nonetheless, these methods remain fundamentally constrained by the limitations of RAG—namely, the inability to fully understand or learn from long-range context that is inaccessible via retrieval alone.
| Category | Dataset | Metrics | AvgL. | Description |
| Accurate Retrieval | SH-Doc QA | Accuracy | 197K | Single-Hop Gold passage retrieval QA. |
| MH-Doc QA | 421K | Multiple-Hop Gold passage retrieval QA. | ||
| LongMemEval (S*) | 355K | Dialogues based QA. | ||
| EventQA | 534K | Novel multiple-choice QA on characters events. | ||
| Test-time Learning | BANKING77 | Accuracy | 103K | Banking intent classification, 77 labels. |
| CLINC150 | Intent classification, 151 labels. | |||
| NLU | Task intent classification, 68 labels. | |||
| TREC Coarse | Question type classification, 6 labels. | |||
| TREC Fine | Question type classification, 50 labels. | |||
| Long Range Understanding | Bench-Sum | F1-Score | 172K | Novel summarization with entity replacement. |
| Detective QA | Accuaracy | 124K | Long-range reasoning QA on detective novels. | |
| Selective Forgetting | FactConsolidation-SH | Accuracy | 262K | Single hop reasoning in facts judgment. |
| FactConsolidation-MH | Multiple hop reasoning in facts judgment. |
@@ 3.1 Dataset Preperation
In this section, we describe how we reconstruct existing datasets and build new ones for evaluating each competency aspect. All datasets with their categories are shown in Table 1. We introduce the details in datasets curation in Appendix A.
@@
We adopt four datasets to evaluate the accurate retrieval capability of memory agents. Three are reconstructed from existing benchmarks, and one is newly created: (1) Document Question Answering: This is a NIAH-style QA task where a long passage contains single (SH-QA) or multiple (MH-QA) documents answering the input question. The agent must identify and extract relevant snippets from the extended context. (2) LongMemEval: This benchmark evaluates memory agents on long dialogue histories. Although task types like information extraction (IE) or multi-session reasoning are included, most tasks can be reformulated as single-retrieval problems requiring agents to retrieve the correct segments spanning a long multi-turn conversation. We reformulated chat history into five long dialogues (355K tokens) with 300 questions (LongMemEval (S*) in Table 1). We create LongMemEval (S*) specifically for increasing the number of questions per context, mitigating the exhaustive needs of reconstructing the memory for each question. (3) EventQA (ours): We introduce EventQA this reasoning style NIAH task to evaluate agents’ ability to recall and reason about temporal sequences in long-form narratives. In this dataset, the agent is required to read a novel and select the correct event from a series of candidates after receiving up-to five previous events. Unlike other long-range narrative text datasets that require extensive manual annotation (Zhang et al., 2024; Xu et al., 2024), our dataset is built through a fully automated pipeline, making the process more efficient and scalable. Moreover, this pipeline can be directly applied to other novel-style texts.
@@
We evaluate TTL via two task categories: (1) Multi-Class Classification (MCC): We reconstructed five classification datasets used in prior TTL work (Bertsch et al., 2024; Yen et al., 2024): BANKING77 (Casanueva et al., 2020), CLINC150 (Larson et al., 2019), TREC-Coarse, TREC-Fine (Li & Roth, 2002), and NLU (Liu et al., 2019). Each task requires the agent to map sentences to class labels, leveraging previously seen labeled examples in context. (2) Recommendation: Based on the setup from (Li et al., 2018; He et al., 2023), we construct a dataset to evaluate movie recommendation via dialogue history. The agent is exposed to thousands of movie-related dialogue turns and is asked to recommend twenty relevant movies based on the long interaction history.
@@
We evaluate LRU via two tasks: (1) Novel Summarization (Summ.): We adopt the Summarization task En.Sum from -Bench (Zhang et al., 2024). The agent is required to analyze and organize the plot and characters of the novel, and then compose a summary of 1000 to 1200 words. (2) Detective QA (Det QA): We also create a difficult question set from Detective QA (Xu et al., 2024), which include ten novels with 71 questions and these questions require agents to do reasoning over a longer narrative range.
Datasets for Selective Forgetting (SF)
To assess whether an agent can forget out of date memory and reason over them, we construct a new dataset called FactConsolidation. Specifically, We build this benchmark using counterfactual edit pairs from MQUAKE (Zhong et al., 2023). Each pair contains a true fact and a rewritten, contradictory version. These are ordered such that the rewritten (new) fact appears after the original, simulating a realistic update scenario. We concatenate multiple such edit pairs to create long contexts of length 6K, 32K, 64K, 262K. We then adpot MQUAKE’s original questions and categorize them into: (1) FactConsolidation-SH (Ours) (SH means Single-Hop), requiring direct factual recall (e.g., “Which country was tool created in?”), and (2) FactConsolidation-MH (Ours) (MH refers to Multi-Hop), requiring inference over multiple facts (e.g., “What is the location of death of the spouse of person ?”). Agents are prompted to prioritize later information in case of conflict and reason based on the final memory state. This setup directly evaluates the strength and consistency of selective forgetting over long sequences.